Two coils A and B are wound on same iron core. There are 500 turns on coil A and 3000 turns on coil B. A current of 5A through coil A produces a flux of 600x10 wb in core. If


this current is reversed in 0.02sec, claculate average e.m.f. Induced in coils A and B.


1
Expert's answer
2021-03-18T03:36:49-0400


In coil B


"E_i=|\frac{\Delta\Phi}{\Delta t}|"


"\Phi=N_A\cdot\Phi_0\to \Phi_0=\Phi/N_A=6000/500=12\ Wb"


"\Phi_B=N_B\cdot \Phi_0=12\cdot3000=36000\ (Wb)"


"E_B=2\cdot\Phi_B/\Delta t=2\cdot36000/0.02=3.6\cdot10^6\ (V)"


In coil A


"E_A=2\cdot\Phi_A/\Delta t=2\cdot6000/0.02=6\cdot10^5\ (V)"